OpenStudy (anonymous):

A racehorse took 2 2/5 minutes to run 1 1/2 miles, how long did it take to run 1 mile?

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@pooja195 HELP ME

OpenStudy (anonymous):

@pooja195 NOW

OpenStudy (anonymous):

@pooja195 PLZ

OpenStudy (anonymous):

@pooja195 DDDDX HEEEELP MEEEEEEEE

TheSmartOne (thesmartone):

Do you know about proportions? Or how to make them, and solve them? (:

OpenStudy (anonymous):

im pretty bad at them T.T

TheSmartOne (thesmartone):

Ok, they have given us three pieces of informations, and they want us to find the fourth piece.

TheSmartOne (thesmartone):

So they gave us this: 2 2/5 minutes to run 1 1/2 miles ????? minutes to run 1 mile

OpenStudy (anonymous):

soo do we divide 2 2/5 by 1 1/2?

TheSmartOne (thesmartone):

Well, first lets simplify 2 2/5 That equals to 7/5 and 1 1/2 = 3/2 So lets rewrite what they gave us so far :) 7/5 minutes to run 3/2miles ????? minutes to run 1 mile

TheSmartOne (thesmartone):

So make a proportion out of that: \(\sf\Huge\frac{\frac{7}{5}~mins}{\frac{3}{2}~mi}=\frac{??~mins}{1~mile}\)

OpenStudy (anonymous):

\[\frac{ 14 }{ 10 }\div \frac{ 15 }{ 10 }?\]

OpenStudy (anonymous):

can u simplify it like that?

TheSmartOne (thesmartone):

yes, you can :)

TheSmartOne (thesmartone):

\[\sf\frac{ 14 }{ 10 }\div \frac{ 15 }{ 10 }=\frac{14}{10}\times\frac{10}{15}\]

OpenStudy (anonymous):

\[\frac{ 140 }{ 150 }\]

TheSmartOne (thesmartone):

And you can remove the extra zeros so you are left with \(\sf\Large\frac{14}{15}\)

TheSmartOne (thesmartone):

So now we have \(\sf\Huge\frac{14}{15}=\frac{??}{1}\)

TheSmartOne (thesmartone):

So do you know how to cross multiply? \(\sf\Large\frac{a}{b} = \frac{c}{d}\to ad=bc\)

OpenStudy (anonymous):

dis is de part where i get all confused ?.?

OpenStudy (anonymous):

oh yeh i do

OpenStudy (anonymous):

14x1=15x

TheSmartOne (thesmartone):

So this is what we have so far: \(\sf\Huge\frac{14}{15}=\frac{??}{1}\) This is what we need to use to solve it: \(\sf\Large\frac{a}{b} = \frac{c}{d}\to ad=bc\)

OpenStudy (anonymous):

so 14=15x?

TheSmartOne (thesmartone):

exactly, and now just divide 15 on both sides! :)

OpenStudy (anonymous):

k so so x=14/15

OpenStudy (anonymous):

and 14/15=0.93... T.T

TheSmartOne (thesmartone):

Correct, but I think they would just want the final answer in fraction. So just 14/15.

OpenStudy (anonymous):

crud, thats not even part of the answers T.T

TheSmartOne (thesmartone):

Oops, I made a mistake :/

TheSmartOne (thesmartone):

\(\color{blue}{\text{Originally Posted by}}\) @TheSmartOne Well, first lets simplify 2 2/5 That equals to 7/5 and 1 1/2 = 3/2 So lets rewrite what they gave us so far :) 7/5 minutes to run 3/2miles ????? minutes to run 1 mile \(\color{blue}{\text{End of Quote}}\) 2 2/5 = 12/5

OpenStudy (anonymous):

O.O THESMARTONE MADE A MISTAKE, IS THE WORLD ENDING!?!

TheSmartOne (thesmartone):

So make a proportion out of that: \(\sf\Huge\frac{\frac{12}{5}~mins}{\frac{3}{2}~mi}=\frac{??~mins}{1~mile}\)

OpenStudy (anonymous):

idk how to make proportions out of fractions T.T (im pretty new to them)

TheSmartOne (thesmartone):

Basically, keep the units the same. So you have minutes over miles on one side. And minutes over miles on the other side :)

OpenStudy (anonymous):

i dont get it, isnt that how it is right now? ?.?

TheSmartOne (thesmartone):

That is how you make i a proportion and that is how it is right now. :) remember it should have the same units on both sides.

OpenStudy (anonymous):

ok so... wut about the fraction parts?

TheSmartOne (thesmartone):

Lets simplify it :) \(\sf\Huge\frac{\frac{12}{5}\times \frac{2}{3}}{\frac{3}{2} \times \frac{2}{3}}=\frac{??~mins}{1~mile}\)

OpenStudy (anonymous):

k gimme a sec to solve this

OpenStudy (anonymous):

1 3/5 over 1?

TheSmartOne (thesmartone):

correct :)

OpenStudy (anonymous):

OH SO THE ANSWER IS 1 3/5 THX!

TheSmartOne (thesmartone):

Correct :D